# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v1

@@ II-C Multi-Static Sensing
$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\boldsymbol{\alpha}$
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\boldsymbol{\mathfrak{h}}+\textbf{n}[m],$ (17)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[25, Sec. V].
II-D Transmit ISAC Precoding Vectors
@@ III Reliability and Delay Analysis for URLLC
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [27, Thm. 6.1] and [25, Lem. 1],
@@ IV Sensing Analysis
We have numerically shown in [19] and [25], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability. When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$. This fact motivates the optimization problems presented in Section VI.
Using the previously defined vector $\boldsymbol{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, from [25], the sensing SINR with independent RCS values is given by
$\mathsf{SINR_{\rm s}}=\frac{\boldsymbol{\rho}^{T}\textbf{A}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}\boldsymbol{\rho}},$ (30)
where
@@
As we see, the sensing SINR is a function of symbols which are random for different blocklength values. When the blocklength is fixed, we can assume that the symbols are known in advance. This implies that although the symbols are known when processing the received sensing signals, they cannot be known during the resource allocation phase. To this end, we take the expectation with respect to the random symbols. Then, the average sensing SINR would be
$\overline{\mathsf{SINR}}_{s}$ $=\frac{\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{A}\}\boldsymbol{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\mathbb{E}\{\textbf{B}\}\boldsymbol{\rho}},$ (33)
The symbols $s_{i}[m]$ are independent, having zero mean and unit variance. Therefore, $\mathbb{E}\{s_{i}^{*}[m]s_{i}[m]\}=1$ and $\mathbb{E}\{s_{i}^{*}[m]s_{j}[m]\}=0$ for $i\neq j$. Then, the expectations in (33) are obtained as follows
$\mathbb{E}\{\textbf{A}\}=L_{d}\,M\textbf{A}_{D},\quad$ $\mathbb{E}\{\textbf{B}\}=L_{d}\,\textbf{B}_{D},$ (34)
where
$\textbf{A}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})$
×$\times\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\Bigg{)}\Bigg{)}$ (35)
$\textbf{B}_{D}=\mathrm{diag}\Bigg{(}\mathrm{diag}\Bigg{(}$ $\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\Bigg{)}\Bigg{)}.$ (36)
The average sensing SINR with respect to symbols and RCS values is finally written as
$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\boldsymbol{\rho}^{T}\textbf{A}_{D}\boldsymbol{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\boldsymbol{\rho}^{T}\textbf{B}_{D}\boldsymbol{\rho}}.$ (37)
@@ V E2E Power Consumption Modeling
Compared to communication networks, ISAC networks are expected to consume more power due to sensing tasks. In general, E2E power consumption in a system with C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [23].
In the considered C-RAN architecture, all processing is done in the cloud. Let $P_{\textrm{AP},0}^{\rm tx}$ and $P_{\textrm{AP},0}^{\rm rx}$ be the static power consumption of the transmit ISAC AP and the receive sensing AP, respectively. The total power consumption, taking into account both communication and sensing, can be expressed as
@@
where $P_{\rm fixed}$ is the fixed power consumption at the cloud which is independent of the load. $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in giga-operations per second (GOPS), respectively[14, 31]. The processing resource utilization can be expressed as
@@ V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [28, App. B], [23].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [28, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\boldsymbol{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. Finally, the multiplications in (11) costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
@@
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [28, App. B] is
@@ V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we compute the GOPS for sensing operations. In downlink, first the sensing precoding vector $\textbf{w}_{0}$ should be obtained. Next operation is multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient which costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. Let $C_{\rm prec-comp}^{\rm s}$ be the number of real multiplications/divisions to compute the sensing precoder $\textbf{w}_{0}$. The unitary matrix U in (19) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [28, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
@@
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. We use the maximum a posteriori ratio test (MAPRT) detector proposed in [25, Lem. 2]. The test statistics is given by
$T$ $=\frac{1}{\sigma_{n}^{2}}\left(\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}\right),$ (54)
where
$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (58)
Let $C_{\rm se-comp}$ denote the number of real multiplications/divisions to compute the test statistics $T$. The computational complexity can be expressed as
$C_{\rm se-comp}=C_{\rm se-prep}+C_{\rm se-detector}$ (59)
where $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are the computational complexity of pre-processing and computing the test statistics for sensing, respectively. The former corresponds to computing (55)-(58), and the latter corresponds to computing the test statistics in (54). The computational complexity for pre-processing part and each step of computing the test statistics are listed in Table I. The total $C_{\rm se-prep}$ and $C_{\rm se-detector}$ are obtained as
$C_{\rm se-prep}=L_{d}C_{{\rm se-prep},m}$
$L_{d}\Bigg{(}20MN_{\rm tx}N_{\rm rx}+16M^{2}N_{\rm tx}N_{\rm rx}+4N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
$\quad+4\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)+8N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}\Bigg{)}$ (60)
$C_{\rm se-detector}=8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
$\quad+8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
$\quad+8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right),$ (61)
where $C_{{\rm se-prep},m}$ is defined as $C_{\rm se-prep}/L_{d}$. Finally, the sensing GOPS is obtained as
@@
$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)}.$ (62)
TABLE I: Number of real multiplications/divisions for computing test statistics.
Operation # of real multiplications/division
$\textbf{G}[m],\quad\forall m$ $20\,L_{d}\,MN_{\rm tx}N_{\rm rx}$
a $8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,$
b $8\,L_{d}\,M^{2}\,N_{\rm tx}\,N_{\rm rx}$
C $4\,L_{d}N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M$
D $4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)$
E $8\,L_{d}\,N_{\rm rx}\,N_{\rm tx}^{2}\,M^{2}$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $8\frac{\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}}{3}$
$\textbf{D}^{-1}$ $8\frac{(M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}}{3}$
$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)$

### v2

@@ II-C Multi-Static Sensing
$\textbf{y}[m]$ $=\underbrace{\mathrm{blkdiag}\left(\textbf{G}_{1}[m],\ldots,\textbf{G}_{N_{\rm rx}}[m]\right)}_{\triangleq\textbf{G}[m]}\bm{\alpha}$
$+\underbrace{\left(\textbf{I}_{N_{\rm rx}}\otimes\left(\textbf{x}^{T}[m]\otimes\textbf{I}_{M}\right)\right)}_{\triangleq\textbf{X}[m]}\bm{\mathfrak{h}}+\textbf{n}[m],$ (16)
where $\textbf{G}_{r}[m]=\begin{bmatrix}\textbf{g}_{r,1}[m]&\cdots&\textbf{g}_{r,N_{\rm tx}}[m]\end{bmatrix}\in\mathbb{C}^{M\times N_{\rm tx}}$, $\textbf{x}[m]=\begin{bmatrix}\textbf{x}^{T}_{1}[m]&\cdots&\textbf{x}^{T}_{N_{\rm tx}}[m]\end{bmatrix}^{T}\in\mathbb{C}^{N_{\rm tx}M}$, and $\mathfrak{h}\sim\mathcal{CN}\left(\textbf{0},\textbf{R}\right)$ is the vectorized target-free channels[34, Sec. V].
II-D Target Detection
We employ two maximum a posteriori ratio test (MAPRT) detectors with two levels of complexity: the clutter-unaware detector [13] and the clutter-aware detector [34, Lem. 2]. The clutter-unaware detector ignores the presence of clutter for the sake of complexity reduction, whereas the clutter-aware detector accounts for unknown clutter. The corresponding test statistics are given by
$T_{\text{c-unaware}}=\textbf{a}^{H}\textbf{C}^{-1}\textbf{a},$ (17)
$T_{\text{c-aware}}=\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}^{H}\left(\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}-\begin{bmatrix}\textbf{0}&\textbf{0}\\ \textbf{0}&\textbf{D}^{-1}\end{bmatrix}\right)\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix},$ (18)
respectively, where
$\textbf{E}=\sum_{m=1}^{L_{d}}\textbf{G}^{H}[m]\textbf{X}[m].$ (22)
II-E Transmit ISAC Precoding Vectors
@@ III Reliability and Delay Analysis for URLLC
Moreover, given that only $\mathbb{E}\left\{\textbf{h}_{i}^{H}\textbf{w}_{i}\right\}$ is known at UE $i$, and according to [35, Thm. 6.1] and [34, Lem. 1],
@@ IV Sensing Analysis
We assess the sensing performance in terms of detection probability under a certain false alarm probability and sensing update rate in terms of the number of sensing measurement updates per time unit. It is numerically shown in [13] and [34], that maximizing the sensing SINR improves the target detection probability under a constant false alarm probability. When studying other sensing tasks, it is naturally desired to keep the sensing SINR above a required threshold denoted by $\gamma_{\rm s}$, i.e., $\mathsf{SINR}_{s}\geq\gamma_{\rm s}$. This fact motivates the optimization problems presented in Section VI. In the following subsections, we present sensing SINR and sensing update rate.
IV-A Sensing SINR
From [34] , the sensing SINR is given by
$\mathsf{SINR_{\rm s}}=\frac{\bm{\rho}^{T}\textbf{A}\bm{\rho}}{L_{d}MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}\bm{\rho}},$ (35)
where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$ and
@@ IV-A Sensing SINR
The sensing SINR is a function of symbols, which vary randomly for different blocklength values. This implies that although the symbols are known when processing the received sensing signals, they cannot be known during the resource allocation phase. To this end, we take the expectation with respect to the random symbols. Thus, the average sensing SINR would be
$\overline{\mathsf{SINR}}_{s}$ $=\frac{M\bm{\rho}^{T}\textbf{A}_{D}\bm{\rho}}{MN_{\rm rx}\sigma_{n}^{2}+\bm{\rho}^{T}\textbf{B}_{D}\bm{\rho}},$ (38)
where $\textbf{A}_{D}$ and $\textbf{B}_{D}$ are diagonal matrices with
$\left[\textbf{A}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\beta_{r,k}\left(\textbf{W}_{k}^{H}\textbf{a}^{*}(\varphi_{k},\vartheta_{k})\textbf{a}^{T}(\varphi_{k},\vartheta_{k})\textbf{W}_{k}\right)_{ii},$ (39)
$\left[\textbf{B}_{D}\right]_{ii}$ $=\sum_{r=1}^{N_{\rm rx}}\sum_{k=1}^{N_{\rm tx}}\mathrm{tr}\left(\textbf{R}_{\mathrm{rx},(r,k)}\right)\left(\textbf{W}_{k}^{H}\textbf{R}^{T}_{{\rm tx},(r,k)}\textbf{W}_{k}\right)_{ii}.$ (40)
IV-B Sensing Update Rate (SUR)
In our system model, the sensing information is updated once throughout each transmission block. Thus, the sensing update rate, denoted as $R_{\rm s}$, is
$R_{\rm s}=\frac{B}{L}\quad(\rm updates/second).$ (41)
@@ V E2E Power Consumption Modeling
Compared to communication networks, ISAC networks are expected to consume more power due to sensing tasks. In general, E2E power consumption in a system with virtualized C-RAN architecture is consisting of two main components: i) the radio site power consumption, including the AP power consumption and ii) the power consumption at the edge cloud, denoted by $P_{\rm cloud}$ [17].
In the considered virtualized C-RAN architecture, all processing is done in the cloud. Let $P_{\textrm{AP},0}^{\rm tx}$ and $P_{\textrm{AP},0}^{\rm rx}$ be the static power consumption of the transmit ISAC AP and the receive sensing AP, respectively. The total power consumption, taking into account both communication and sensing, can be expressed as
@@
where $P_{\rm fixed}$ is the fixed power consumption at the cloud which is independent of the load. $P_{\rm cloud,0}^{\rm proc}$ is the processing power consumption in the idle mode. $\sigma_{\rm cool}\in(0,1]$ and $\Delta_{\rm cloud}^{\rm proc}$ denote the cooling efficiency of the cloud and the slope of the load-dependent power consumption for processing at the digital unit (DU) in the cloud, respectively. Moreover, $C_{\rm max}$ and $C_{\rm cloud}\in[0,C_{\rm max}]$ are the maximum processing capacity of the processing resources and the total processing resource utilization in GOPS, respectively[30, 39]. The processing resource utilization can be expressed as
@@ V-A GOPS Analysis of Digital Operations for Communication at the Cloud
In this section, we analyze the GOPS for digital signal processing corresponding to the communication tasks including the uplink channel estimation and downlink transmission. To compute the number of real multiplications, we mainly follow the GOPS analysis in [36, App. B], [17].
Let $C_{\textrm{ch-est}}$ denote the computational complexity of the LMMSE channel estimation approach for all the APs. To compute the channel estimates, we first obtain the vectors $\textbf{y}_{t_{i},k}^{p}$. From [36, App. B], the multiplication of one matrix of size $M\times L_{p}$ with a vector of size $L_{p}\times 1$ results in $ML_{p}$ complex multiplications. Hence, obtaining $\textbf{y}_{t_{i},k}^{p}$ for $N_{\rm ue}$ UEs at all $N_{\rm tx}$ APs costs $8ML_{p}N_{\rm ue}N_{\rm tx}$ real multiplications/division in total, if $L_{p}\geq N_{\rm ue}$. Otherwise, the number of real multiplications/divisions would be $8ML_{p}^{2}N_{\rm tx}$. Moreover, we need to compute the matrices $\textbf{R}_{i,k}^{\prime}$ and $\bm{\Psi}_{t_{i},k}^{\prime}$, given that $\bar{\textbf{h}}_{i,k}$ and $\textbf{R}_{i,k}$ are known. However, this pre-computation can be neglected since the channel statistics are usually constant for a while and there is no need to compute them every coherence block. Finally, the multiplications in (10) costs $8M^{2}N_{\rm ue}N_{\rm tx}$ real operations. Then, $C_{\textrm{ch-est}}$ is equal to
@@
The number of real multiplications/divisions to compute centralized RZF precoding vector for all the UEs from [36, App. B] is
@@ V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we present the GOPS analysis for sensing operations, including signal transmission and processing the received signal. The GOPS corresponding to the sensing transmissions consists of computing the sensing precoding vector $\textbf{w}_{0}$, denoted by $C_{\rm prec-comp}^{\rm s}$, and obtaining the sensing signal by multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient. Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. The unitary matrix U in (24) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [36, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
@@
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. Hence, the sensing GOPS is obtained as
@@
$C_{\rm proc}^{\rm s}$ $=\frac{B}{L10^{9}}\Bigg{(}\underbrace{12L_{d}MN_{\rm tx}}_{\textrm{precoding + power}}+C_{\textrm{prec-comp}}^{\rm s}+\underbrace{C_{\rm se-comp}}_{\textrm{target detection}}\Bigg{)},$ (49)
where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics. Given the clutter-unaware detector in (17), the computational complexity is computed as
$C_{\rm se-comp}^{\rm c-unaware}$ $=L_{d}\Bigg{(}\underbrace{20MN_{\rm tx}N_{\rm rx}}_{\rm{for}\,\textbf{G}}+\underbrace{8\,M^{2}\,N_{\rm rx}\,N_{\rm tx}}_{\rm{for}\,\textbf{a}}$
$\quad+\underbrace{4\,N_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)M}_{\rm{for}\,\textbf{C}}\Bigg{)}+\underbrace{8\frac{(N_{\rm tx}^{3}-N_{\rm tx})N_{\rm rx}}{3}}_{\textrm{for inverse {C}}}$
$\quad+\underbrace{8\left((N_{\rm tx}N_{\rm rx})^{2}+N_{\rm tx}N_{\rm rx}\right)}_{\textrm{for final step}},$ (50)
where we used the block diagonal structure of C in getting the inverse of it.
The computational complexity for each step of computing the test statistics of clutter-aware detector in (18) are listed in Table I.
TABLE I: Computational complexity with clutter-aware MAPRT detector
Operation Computational complexity
$\textbf{G}[m],\quad\forall m$ $\mathcal{O}\left(20\,L_{d}\,MN_{\rm tx}N_{\rm rx}\right)$
a $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\right)\,$
b $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,\right)$
C $\mathcal{O}\left(4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)\right)$
D $\mathcal{O}\left(4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)\right)$
E $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}\right)$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
$\textbf{D}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)\right)$
$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $\mathcal{O}\left(8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$